# Shuffle-R1: Efficient RL framework for Multimodal Large Language Models via Data-centric Dynamic Shuffle

### v3

@@ 6 Conclusion
In this paper, we propose Shuffle-R1, a simple but effective framework that improves the training efficiency of reinforcement learning of multimodal large language models. Through Pairwise Trajectory Sampling and Advantage-based Batch Shuffle, our framework significantly outperforms representative algorithms and models in both in-domain and out-of-domain tasks, demonstrating the value of data-centric adaptive design. We hope that our motivations, method, and findings are helpful for further research.
@@ Appendix B Prompt Design
Appendix C Experiment Settings
@@
Reward Calculation.
[…]
The final reward is the weighted sum of above rewards:
Hyperparameters.
We use EasyR1 (Yaowei et al. 2025) as our training framework. Full hyperparameter settings during training is shown in Tab. 6. For experiments on Geometry3K and K12 with $\sim 2.1k$ training samples, we set the training steps to 80. For the joint training experiments ($\sim 30k$ training samples), we increase the training steps to 150 due to extended data size. Other hyperparameters that are not mentioned are kept to default values of EasyR1.
@@
Table 6: Hyperparameter settings.
@@
All experiments are conducted on $8\times$ 80G-SXM GPUs.
Appendix D Dataset Collection
For training dataset, we choose Geometry3K (Lu et al. 2021) and MMK12 dataset. Geometry3K dataset is a high quality real-world geometry problem solving dataset, containing 2.1K training samples and 601 test samples. It contains a wide range of geometry problems with varying levels of difficulty. The text problems are very compact with most of the geometric conditions represented in images, making it suitable for our training. MMK12 dataset is introduced by MM-Eureka (Meng et al. 2025), which contains 16k math reasoning training samples. The training samples have both geometric and non-geometric problems and have been carefully examined and manually filtered to ensure quality. The test set of MMK12 further includes other STEM problems such as physics, biology and chemistry. During our training, we used a randomly selected subset of MMK12 provided by NoisyRollout (Liu et al. 2025a) (referred to as ‘K12’ in the paper to distinguish it from full set of MMK12), which has 2.1k samples with the same size as Geometry3K. All the training samples are in free-form format. Examples of training data shown in Fig. 7.
@@ Appendix D Dataset Collection
To fully evaluate model’s reasoning ability, apart from in-domain test set from training dataset, we select several representative benchmarks to examine model’s performance on math reasoning, visual perception, and chart understanding tasks. For math reasoning task, we choose MathVerse (Zhang et al. 2024b), MathVision (Wang et al. 2024), WeMath (Qiao et al. 2024) and MathVista (Lu et al. 2023). We choose HallusionBench (Guan et al. 2024) and ChartQA (Masry et al. 2022) for visual perception and chart understanding tasks respectively. We believe that a good RL algorithm can not only improve in-domain performance, but also have the potential to promote robust generalization to out-of-domain tasks. Examples of evaluation data shown in Fig. 8.
@@
Appendix E Evaluation Metrics
@@
For evaluation of MathVerse, MathVision, WeMath, MathVista and ChartQA, we employ Gemini-2.0-Flash-001 (Deepmind 2025) to first extract the predicted answer from model response then compare it with ground truth. Fig. 10 and Fig. 10 demonstrate the extraction and verification prompt for Gemini. Specifically, we report accuracy on WeMath under loose mode, and overall accuracy on MathVerse (including all sub categories: Text Dominant, Text Lite, Vision Intensive, Vision Dominant and Vision Only).
@@
Appendix F More Experimental Results
Detailed Model Performance.
Comparison with more RL algorithms.
We provide additional comparisons with more representative RL algorithms, i.e. RLOO (Ahmadian et al. 2024) and REINFORCE++ (Hu 2025). We conduct experiments on Qwen2.5-VL-3B-Instruct with Geometry3K dataset. All the experiment settings are kept the same as GRPO and DAPO in main paper. The final model performance is shown in Tab. 8. Our framework outperform these algorithms by a large margin in both in-domain and out-of-domain tasks.
Qwen2.5-VL-3B 34.77 21.94 58.40 51.72 59.83 73.08 49.96
+ Ours (Geo3K) 43.55 25.30 61.80 64.14 63.19 77.04 55.84
+ Ours (K12) 44.06 26.48 64.90 64.77 65.72 78.28 57.37
Qwen2.5-VL-7B 42.59 25.76 67.40 63.51 65.19 79.84 57.38
+ Ours (Geo3K) 50.96 27.47 70.90 69.19 69.51 81.64 61.61
+ Ours (K12) 48.59 28.61 73.20 69.71 69.87 82.60 62.09
Qwen2.5-VL-3B 25.79 34.77 21.94 58.40 51.72 59.83 73.08
+ RLOO 42.09 39.94 22.96 58.90 59.48 64.24 76.68
+ REINFORCE++ 41.76 41.70 24.86 60.90 63.51 62.99 76.20
+ Ours 47.88 43.55 25.30 61.80 64.14 63.19 77.04
Comparison with Direct Experience Replay.

### v4

@@ 5 Conclusion
In this paper, we introduce Shuffle-R1, a simple but effective framework designed to improve the training efficiency of reinforcement learning of multimodal large language models. Through Pairwise Trajectory Sampling and Advantage-based Batch Shuffle, our framework significantly outperforms representative algorithms and models in both in-domain and out-of-domain tasks, demonstrating the value of data-centric adaptive design. We hope that our motivations, method, and findings are helpful for further research.
Acknowledgment
This work was supported by the National Natural Science Foundation of China (NSFC) under Grants No. 62225603 and 623B2038.
Ethics statement
All training and evaluation data used in this project are collected from public academic datasets. All the pre-trained model checkpoints and code framework employed are forked from open-source research projects. During the research process, we strictly adhered to the open-source licenses and usage requirements of all relevant parties. No non-public / private data was used.
Furthermore, all authors strictly complied with the ICLR Code of Ethics throughout the research process and affirm that no activities violating these ethical standards were involved.
Reproducibility statement
We have reported the base model checkpoints, code framework, training data, evaluation benchmarks and metrics, and core hyperparameter settings used in the project in the main paper. A detailed pseudo code explaining our framework flow is provided in Appendix A. More detailed hyperparameter settings and information about the training infrastructure are provided in Appendix D. In addition, we have also included the prompts used during training and inference, examples of the training and evaluation data, as well as examples of the model inference results in Appendix EF and G.
LLM Usage Statement
During this research process, the usage of LLM is strictly restricted and was used in limited ways: (1) GitHub Copilot was employed during code development for code checking and bug fixes; (2) LLM-based tools were used during paper writing for language refinement and grammar correction, and (3) LLM judges were applied as a part of evaluation metrics, including key answer extraction and correctness assessment.
All core components of this work, including research motivation, algorithm design, core code implementation, data collection, and experiment analysis, and experimental conclusions, were conducted entirely and independently by the authors without reliance on LLMs.
@@ Appendix B Prompt Design
Appendix C More Experimental Results
Detailed Model Performance.
Full Setting Comparison.
Improved training algorithm leads to improved performance. To better reveal the model performance beyond the limited data scale of Geo3K, we conduct full 30k-sample 150-step experiment on Qwen2.5-VL-7B. As shown in Tab. 10, Shuffle-R1 achieves clear and consistent gains over strong baselines across all six evaluation benchmarks. This result further proven the robustness and scalability of Shuffle-R1.
Comparison with more RL algorithms.
We provide additional comparisons with more representative RL algorithms, i.e. RLOO (Ahmadian et al., 2024) and REINFORCE++ (Hu, 2025). We conduct experiments on Qwen2.5-VL-3B-Instruct with Geometry3K dataset. All the experiment settings are kept the same as GRPO/DAPO/GSPO in main paper. The final model performance is shown in Tab. 11. Our framework outperform these algorithms by a large margin in both in-domain and out-of-domain tasks.
Qwen2.5-VL-3B 34.77 21.94 58.40 51.72 59.83 73.08 49.96
[…]
Qwen2.5-VL-7B 42.59 25.76 67.40 63.51 65.19 79.84 57.38
[…]
Table 10: Full 30k experiment on Qwen2.5-VL-7B. Highest accuracy marked in Bold.
Qwen2.5-VL-7B 42.6 25.8 67.4 63.5 65.2 79.8 57.4
+ GRPO 50.6 28.3 74.5 69.7 70.7 81.4 62.5
+ DAPO 51.4 28.8 75.3 71.5 71.0 82.8 63.4
+ GSPO 50.8 28.2 75.3 70.1 69.7 82.9 62.8
[…]
Qwen2.5-VL-3B 25.79 34.77 21.94 58.40 51.72 59.83 73.08
+ RLOO 42.09 39.94 22.96 58.90 59.48 64.24 76.68
+ REINFORCE++ 41.76 41.70 24.86 60.90 63.51 62.99 76.20
[…]
Comparison with Prioritize Experience Replay.
Detailed Impact of Hyperparameters.
Here, we provide detailed model performance comparison under different hyperparameter settings in Fig. 12, as supplementary information to corresponding ablation experiments in main paper. For sampling ratio ($\alpha$), both the in-domain and out-of-domain accuracy improves as we increase $\alpha$ from 0.25 to 0.5. The model performance shows a mild decrease when $\alpha$ is further raised to 0.75. This result demonstrates again that not all the rollouts contribute equally positive to RL training. For shuffle times ($S$), the model performance receives a consistent gain when $S$ is increased from 4 to 8, but declines when $S$ is set to 16. This result suggests that appropriate repeated training on high-quality samples can enhance the model’s reasoning ability without affecting data diversity.
Table 12: Performance of Shuffle-R1 under different hyperparameter settings. Highest accuracy marked in Bold.
Setting Geo3K MathVerse MathVision MathVista WeMath HallBench ChartQA
Ablation on sampling ratio $\alpha$ ($S=8$)
$\alpha=0.25$ 45.75 42.43 23.05 60.20 63.62 63.09 75.36
[…]
$\alpha=0.75$ 47.41 41.70 25.62 60.90 64.54 62.99 73.36
Ablation on shuffle times $S$ ($\alpha=0.5$)
$S=4$ 45.92 42.21 24.78 62.30 64.08 62.25 75.60
[…]
$S=16$ 47.23 43.70 24.86 60.90 62.64 62.20 74.84
Appendix D Experiment Settings
@@
Table 13: Hyperparameter settings.
@@
Reward Calculation.
[…]
The final reward is the weighted sum of above rewards:
Hyperparameters.
We use EasyR1 (Yaowei et al., 2025) as our training framework. Full hyperparameter settings during training is shown in Tab. 13. For experiments on Geometry3K and K12 with $\sim 2.1k$ training samples, we set the training steps to 80. For the joint training experiments ($\sim 30k$ training samples), we increase the training steps to 150 due to extended data size. Other hyperparameters that are not mentioned are kept to default values of EasyR1.
@@
All experiments are conducted on $8\times$ NVIDIA H800-80G-SXM GPUs.
Appendix E Dataset Collection
@@
For training dataset, we choose Geometry3K (Lu et al., 2021) and MMK12 dataset. Geometry3K dataset is a high quality real-world geometry problem solving dataset, containing 2.1K training samples and 601 test samples. It contains a wide range of geometry problems with varying levels of difficulty. The text problems are very compact with most of the geometric conditions represented in images, making it suitable for our training. MMK12 dataset is introduced by MM-Eureka (Meng et al., 2025), which contains 16k math reasoning training samples. The training samples have both geometric and non-geometric problems and have been carefully examined and manually filtered to ensure quality. The test set of MMK12 further includes other STEM problems such as physics, biology and chemistry. During our training, we used a randomly selected subset of MMK12 provided by NoisyRollout (Liu et al., 2025a) (referred to as ‘K12’ in the paper to distinguish it from full set of MMK12), which has 2.1k samples with the same size as Geometry3K. All the training samples are in free-form format. Examples of training data shown in Fig. 11.
To fully evaluate model’s reasoning ability, apart from in-domain test set from training dataset, we select several representative benchmarks to examine model’s performance on math reasoning, visual perception, and chart understanding tasks. For math reasoning task, we choose MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024), WeMath (Qiao et al., 2024) and MathVista (Lu et al., 2023). We choose HallusionBench (Guan et al., 2024) and ChartQA (Masry et al., 2022) for visual perception and chart understanding tasks respectively. We believe that a good RL algorithm can not only improve in-domain performance, but also have the potential to promote robust generalization to out-of-domain tasks. Examples of evaluation data shown in Fig. 11.
Appendix F Evaluation Metrics
@@
For evaluation of MathVerse, MathVision, WeMath, MathVista and ChartQA, we employ Gemini-2.0-Flash-001 (Deepmind, 2025) to first extract the predicted answer from model response then compare it with ground truth. Fig. 13 and Fig. 13 demonstrate the extraction and verification prompt for Gemini. Specifically, we report accuracy on WeMath under loose mode, and overall accuracy on MathVerse (including all sub categories: Text Dominant, Text Lite, Vision Intensive, Vision Dominant and Vision Only).
@@ Appendix G Case Study
Appendix H Theoretical Analysis of Shuffle-R1
In this section, we provide an intuitive analysis on how the adaptive selection & resampling improve the training dynamics, and discuss the bias introduced by this process. We begin by assuming the Advantage($A$) follows a Gaussian distribution, $A\sim\mathcal{N}(0,\sigma^{2})$, which matches the observation of our probe analysis.
We define the original expectation of $|A|$ as:
$E_{\text{orig}}=\mathbb{E}[|A|].$ (13)
We simplify the PTS as a truncate process with a threshold $\tau>0$, and define the kept set $S^{\text{PTS}}=\{|A|>\tau\}$, the corresponding conditional expectation is:
$E_{\text{PTS}}=\mathbb{E}[|A|~|~|A|\geq\tau].$ (14)
On the truncated set, we sample with weights $w(a)\propto|a|$, and the resulting expectation is denoted as:
$E_{\text{new}}=\mathbb{E}[w(a)|a|~|~|a|\in S^{\text{PTS}}].$ (15)
H.1 Intuitive Analysis on Training Dynamics
Proposition 1 (Amplification on expectation of Advantage).
Proof. Define $X=|A|$ on the $S^{\text{PTS}}$, the weighted sampling probability density can be expressed as:
$p_{\text{new}}(x)=\frac{x\cdot p(x~|~S^{\text{PTS}})}{\mathbb{E}[X~|~S^{\text{PTS}}]}=\frac{x\cdot p(x~|~S^{\text{PTS}})}{E_{\text{PTS}}}.$ (16)
Thus the weighted sampling expectation can be calculated as:
$E_{\text{new}}=\int x\cdot\frac{x\cdot p(x~|~S^{\text{PTS}})}{E_{\text{PTS}}}dx=\frac{\mathbb{E}[X^{2}|S^{\text{PTS}}]}{\mathbb{E}[X|S^{\text{PTS}}]}=\frac{E_{\text{PTS}}^{(2)}}{E_{\text{PTS}}}.$ (17)
Based on Cauchy–Schwarz inequality, we can derive that:
$\mathbb{E}[X^{2}]\geq(\mathbb{E}[X])^{2}\Longrightarrow E_{\text{new}}\geq E_{\text{PTS}}$ (18)
Further, the original expectation can be decomposed as:
$E_{\text{orig}}=(1-p_{\tau})\mathbb{E}[|A|~|~|A|<\tau]+p_{\tau}\mathbb{E}[|A|~|~|A|\geq\tau],\quad\text{where}\ p_{\tau}=\mathbb{P}(|A|\geq\tau).$ (19)
Since
$\mathbb{E}[|A|~|~|A|<\tau]<\tau<\mathbb{E}[|A|~|~|A|\geq\tau].$ (20)
we have $E_{\text{PTS}}>E_{\text{orig}}\Longrightarrow E_{\text{new}}>E_{\text{orig}}$. $\qquad\square$
Proposition 2 (Amplification on gradient norm).
Proof. In the policy gradient training, the per-sample gradient can be expressed as:
$g=A\cdot u.\qquad u=\nabla_{\theta}\text{log}\pi_{\theta}(o).$ (21)
g=|A|u.||g||=|A|~||u||. (22)
Derived from the standard assumptions of policy gradient that $u||$ is approximately independent from $|A|$, we have:
[…]
which in turn yields larger update magnitudes per optimization step. $\qquad\square$
H.2 Bias Under Selective Sampling
Proposition 3 (Adaptive sampling introduce positive bias).
Proof. Under the Gaussian distribution assumption, the expectation can be calculated as:
$\mathbb{E}_{\text{new}}[A]$ $=\frac{1}{Z}(\int_{-\infty}^{-\tau}a|a|p(a)da+\int_{\tau}^{\infty}a|a|p(a)da)$
$=\frac{2}{Z}\int_{\tau}^{\infty}a^{2}p(a)da>0$
where the normalized factor $Z$ is defined as:
$Z$ $=\int_{-\infty}^{\infty}|a|p(a)\mathbb{I}(|a|>\tau)da$
$=2\int_{\tau}^{\infty}a\cdot p(a)da$
The positive bias indicates that the gradient update is systematically tend to trajectories that has higher advantage. $\qquad\square$